Let "A = \begin{pmatrix}
 1 & 0 & 3\\
 0 & 4 & 5 \\
 1 & 2 & 6
\end{pmatrix}"

What is the contactor of the entry A23 = 5


  1. 2
  2. -2
  3. 10
  4. -10
1
Expert's answer
2021-07-04T17:23:19-0400

The co-factor of 5 is obtained by eliminating the row and column of 5. By doing this, we obtain the matrix below

"\begin{pmatrix} 
1 & 0 \\ 1&2
\end{pmatrix}"

the determinant of the matrix above is "1(2) -0=2"

Next, we check the sign associated with matrix. The matrix sign used to write the co-factor of each elements is given below;

"\begin{pmatrix} 
+ & - & +\\ -&+&-\\+&-&+
\end{pmatrix}"

Therefore the sign associated with 5 is -1. Hence the co-factor of 5 is "-1(2) = -2"